a rectangle pool is 34' x18'. around pool is rectangluar walkway 50' by 30'. Find the number of square feet of the walkway around the pool?
Answer by checkley77(12569)

34*18=612 ft^2
50*30=1,500 ft^2
1,500-612=888 ft^2 is the area of the walkway.